OpenStudy (anonymous):

solve the system by substitution. 2x+y=-11 3x-4y=11 Help please! will give medal

OpenStudy (adajiamcneal):

ok so looking at the first one. basically you have to multiply one or both of the equations by a number so that the coefficient of one of the variables is the same so that when you subtract/add them then it will cancel. ^that sounds confusing, here's an ex for elimination: x+y=5, 2x+y = 30 so if we multiplied the first equation by 2, we'd get: 2x+2y=10 right? so we can subtract now... 2x+2y=10 - 2x+y=30 ------------ 0 + y = -20 ------------- and now we can plug in -20 for y back in the equation to find x. hope that made sense

OpenStudy (anonymous):

thank you